Determine whether the following function are suitable as wavefunctions:
F(x) = 0 for x < 1;
= 1/x for x > 1;
= 0 for x > 1000.
1
Expert's answer
2020-09-29T06:29:01-0400

For the following basis of functions (Y2p-1, Y2p0 , and Y2p+1), construct the matrix representation of the Lx operator (use the ladder operator representation of Lx). Verify that the matrix is hermitian. Find the eigenvalues and corresponding eigenvectors. Normalize the eigenfunctions and verify that they are orthogonal. Y2p-1 = 1 8p1/2 è æ ø Zö a 5/2 re-zr/2a Sinq e-if Y2po = 1 p1/2 è æ ø Zö 2a 5/2 re-zr/2a Cosq Y2p1 = 1 8p1/2 è æ ø Zö a 5/2 re-zr/2a Sinq eif

.